# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 2.1.1 Instruction Following Tasks in General
In addition to task-oriented evaluations, a complementary line of work has explored abstract reasoning benchmarks designed to test LLMs’ core reasoning capabilities in multi-turn settings. AQA-Bench [29] evaluates LLMs in interactive environments requiring sequential decision-making, such as binary search and graph traversal (DFS, BFS). It emphasizes memory maintenance, procedural adherence, and planning over multiple turns, with both algorithmic and “embodied” (narrative) settings. WILT (Wason Inductive Logic Test) [35] by Banatt et al. (2024) further assesses inductive reasoning by asking models to iteratively discover hidden rules through evidence gathering and hypothesis testing. These benchmarks are task-agnostic and explicitly designed to avoid memorization, targeting core skills such as logical consistency, strategic exploration, and hypothesis refinement.
@@
Besides these abstract reasoning evaluations, other studies have addressed multi-turn interactions from specialized task perspectives. For instance, WebLINX [28] tackles the problem of conversational web navigation, where an LLM-based agent must follow user instructions via dialogue to accomplish tasks on real websites. Tasks range from booking tickets to finding information, requiring the agent to understand natural language commands and manipulate a web page accordingly (click links, fill forms, etc.) over multiple turns. MULTITURNINSTRUCT [45] proposed a systematic benchmark to probe LLMs’ ability to handle sequential, potentially conflicting instructions in a conversation. SysBench [33] is a benchmark specifically designed to evaluate how well LLMs adhere to system-level instructions (the hidden directives guiding an AI assistant) in multi-turn interactions. Emphasizing clear and explicit user intentions, we categorize two recent recommendation system works in this subsection: SAPIENT [46] and ECR [47]. SAPIENT [46] is a multi-turn conversational recommender system that integrates a planning module for strategic dialogue management, while ECR [47] introduces an “empathetic”conversational recommender that enhances traditional recommendation dialogues with emotional awareness.
@@ 2.1.2 Instruction Following Tasks in Math
LLMs have demonstrated impressive performance in solving math problems via single-turn prompts, often generating detailed, step-by-step “chain-of-thought” solutions. For example, providing a few worked examples allows a 540B-parameter model to achieve state-of-the-art accuracy on the GSM8K math word problems [48]. However, complex mathematical tasks frequently require LLMs to engage in multi-turn, instruction-following interactions that involve incremental reasoning, clarification questions, and iterative refinement based on interactive feedback [31, 49]. In these instruction-following math tasks, users iteratively guide LLMs by providing clarifications, corrections, or additional contextual instructions. Such dynamic interactions not only enhance the models’ reasoning processes but also facilitate effective use of external computational tools, enabling the models to perform sophisticated tasks like simulating scenarios or executing complex calculations through code [49, 50]. Instruction-following tasks in math also encompass advanced skills such as follow-up questioning, errors diagnosing and correcting, and educational feedback delivering, thus capturing broader capabilities essential for deploying LLMs in diverse educational and problem-solving contexts.
Several approaches have leveraged multi-turn dialogue with LLMs to enhance mathematical reasoning. For example, Wu et al. (2024) [20] introduced MathChat-Agent, a framework where an LLM collaborates with a user-proxy agent (responsible for tool use, such as a Python solver) via iterative conversation. This approach solved competition-level math problems more effectively, improving accuracy by roughly 6% over standard single-turn chain-of-thought prompts. Similarly, Keating (2024) [51] employed a multi-agent strategy in which two GPT-4 agents engage in debate-style interactions to reach a solution. This dual-agent zero-shot method achieved about 62.7% accuracy on the MATH benchmark—surpassing single-agent baselines and illustrating the benefits of peer deliberation in reasoning. Moving further, Xiong et al. (2024) [52] propose a method to train LLMs to better combine tool usage with their own reasoning for complex tasks. It gathers trajectory-level feedback from users over multiple turns, to refine problem-solving strategies. The learning process is modeled as a Markov decision process (MDP) and adapts direct preference learning algorithms to multi-turn interactions that include external messages. In practice, the model is trained on multi-turn chats where a user first asks a question and then gives Python outputs in later turns. The accuracy of the method is evaluated on the GSM8K and MATH test sets.
@@
Several benchmarks now target multi-turn math dialogues. MathChat-Bench [31] extends GSM8K with four new tasks: follow-up QA, problem generation, error correction, and error analysis. The original problems are modified using GPT-4 to meet specific requirements. For instance, in the follow-up QA task, three rounds of dialogue are created: first, GSM8K test problems with ground-truth answers are presented; then GPT-4o generates two follow-up questions; finally, GPT-4 produces the final answers, which are verified or revised by two other LLMs and human annotators. LLMs are evaluated by both accuracy and scores, assigned by GPT-4, which measure their instruction follow-up abilities. Evaluations of state-of-the-art models on MathChat showed that while they excel at standard one-shot math questions, performance drops sharply on these multi-turn interactions requiring sustained reasoning and dialogue comprehension. To address this gap, the authors also released MathChatSync, a synthetic dialogue dataset for fine-tuning LLMs on conversational math problem solving. Fine-tuning on MathChatSync yielded notable improvements in multi-turn performance.
More broadly, the MINT benchmark [23] evaluates multi-turn tool use and user feedback across domains (including math reasoning). MINT provides an automated framework where the LLM can call a Python interpreter and receive natural language feedback (simulated by GPT-4) in successive turns. Findings from MINT show that multi-turn tool-aided dialogues consistently improve problem-solving success (each tool use or feedback turn yields additional 1–17% accuracy gains). Interestingly, this evaluation also found that some models fine-tuned only on single-turn instructions (via standard supervised tuning or RLHF) underperform in multi-turn settings, suggesting that multi-turn-specific training is needed to excel in interactive math tasks.
Noticeably, several studies have found that LLMs struggle with generalizing to new problems. For example, Liang et al. (2024) [31] show that math-specific LLMs lack adaptive behavior, and their difficulty with generating novel problems highlights their rigidity. Similarly, Macina et al. (2023) [53] report that dialogue tutoring models do not generalize well to unseen math problems. To improve multi-turn math problem solving and generalization, many researchers propose using SFT [31, 53, 23] or RLHF [23, 52]. Although RLHF can affect LLM–tool interactions and leveraging feedbacks [23], studies consistently find that fine-tuning with preference data and instructions boosts downstream performance. Macina et al. (2023) [53] demonstrate that small, fine-tuned models perform significantly better, in terms of correctness and equitable tutoring, than prompting a large model like ChatGPT.
@@ 2.1.3 Instruction Following Tasks in Coding
LLMs often struggle to produce correct code and perform self-debugging in a single pass, frequently requiring multi-turn or iterative interactions [54, 55, 56]. Instruction-following tasks in coding contexts commonly involve collaborative, iterative interactions, where users provide detailed instructions that the LLM translates into executable code. Through subsequent turns, the LLM iteratively integrates feedback, refines initial solutions, strategically plans modifications, executes and tests submodules, and performs debugging until satisfying the given specifications. Such iterative instruction-following tasks are essential to realistically evaluate LLM performance and robustness in dynamic coding scenarios.
@@
Different frameworks and playgrounds are developed for evaluaing code generation qualities [22, 57]. InterCode [22] introduces a generation pipeline for evaluating LLM coding quality through multi-turn interactions that simulate a real-world coding environment. The InterCode framework requires as input a natural language prompt paired with either an answer or a correct code block. The LLM is evaluated using one of three strategies: “Try-again”, where the execution output is fed back as an observation; ReACT, which terminates once the thought chain is complete; or Plan & Solve, which terminates when the plan is fully executed. The experiments involve three programming environments with Spider [58], MBPP [59] and NL2Bash [60] datasets. Zheng et al. (2024) [57] propose a framework for systematically evaluating various prompting techniques for multi-turn code generation by LLMs. Their evaluation is conducted in a zero-shot setting using two competitive coding benchmarks, CodeContests [61] and TACO [62]. PyBench [63] proposes a unified benchmark to evaluate LLM Python coding ability in several categories such as chart analysis, software development, etc. For each task, LLMs are interacting with a code interpreter for a few turns before making a formal response. The LLMs are evaluated on the success rates and average turns to complete a unit test for each task.
@@
Toward effective code generations, CodeGEN [18] and CodeGEN2 [64] are a family of LLMs designed to generate programs from natural language descriptions in multiple turns. The models are evaluated using the Pass Rate metric on their Multi-Turn Programming Benchmark (MTPB), which comprises 115 expert-written problems. Each problem includes multi-step natural language prompts, created by human annotators who decompose the problem into sequential steps. Models are required to generate the complete solution from scratch. Chen et al. (2025) [65] propose to fine-tune existing LLMs with SFT and DPO. They introduce CodeSteer, a framework that guides LLMs through multiple rounds of interaction to generate code. In this system, the primary model (TaskLLM) produces responses, both in natural language and in code, while a supervisory agent (CodeSteerLLM) reviews these outputs using symbolic reasoning and self-answer checking to ensure correctness and provide refined guidance. They are fine-tuned and evaluated on subsets of SymBench, which comprises 37 symbolic tasks with adjustable complexity and includes a synthesized dataset of multi-round guidance/generation trajectories and guidance comparison pairs. OpenCodeInterpreter [66] proposes to fine-tune LLMs with CodeFeedback, a dataset of challenging LeetCode questions, incorporating multi-turn execution feedback (with code interpreters) and dialogues of human (synthesized by GPT-4). CodeAct [67] collect an instruction-tuning dataset, CodeActInstruct, which contains 7,000 multi-turn interactions. PyInstruct [63] is used in PyBench for continuous pretraining and fine-tuning. Zheng et al. (2024) [57] explores a wide range of prompting strategies for effective code generation, focusing on automatic re-prompting over multiple runs.
@@
SQL generation is a subcategory of coding generation that focuses more on data acquisition through large-scale data warehouses. Two benchmarks for SQL generation multi-turn LLMs are identified during the literature search. MMSQL [38] focuses on text-to-SQL generation tasks and introduces a Multi-type and Multi-turn text to-SQL test suite, which is a comprehensive benchmark engineered to evaluate the proficiency of LLMs in handling multi-turn text-to-SQL tasks across diverse question types. MMSQL contains a multi-agent framework anchored by a core Question Detector and Question Decomposer tasked with identifying question types and determining appropriate answering strategies. The framework includes two supportive agents: the Schema Selector, which identifies and provides the essential subset of a database schema, and the SQL Refiner, which is dedicated to refining SQL queries. EHRAgent [26] demonstrates a specific application of SQL generation in healthcare settings, which speeds up the extraction and interaction of clinician information within electronic health record (EHR) systems. EHRAgent [26] translates EHR question-answering into a tool-use planning process, which integrates query-specific medical information and formulates executable code plans through multi-turn dialogues. The model is evaluated based on their ability to reason across multiple tables and generate accurate, actionable insights from complex EHR data.
@@
For education applications, TreeInstruct [68] transforms LLMs into instructor agents that guide users in debugging and writing better code. Acting as a Socratic educator, it plays two roles: the instructor generates tree-based sequential questions, and the verifier identifies tasks to help students understand, assess, and correct their code. To evaluate TreeInstruct, the authors created MULTIDEBUG, a dataset derived from popular LeetCode problems with expert-injected syntactical and conceptual bugs. Evaluation measures include both qualitative factors (relevance, indirectness, logical flow) and quantitative metrics (success rate).
Several compelling research questions arise as common themes in adapting LLMs for effective code generation through multi-turn interactions. For instance, how can an LLM decide whether to employ textual reasoning or programmatic solutions when explicit cues are absent [69, 55]? What iterative interaction protocols enable the model to refine its solution before submitting a final answer [22, 57, 18, 64, 67]? And how should we evaluate LLMs’ coding performance across different programming languages and tasks in diverse domains [22, 63]?
@@ 2.1.4 Discussions
The evaluation landscape has diversified considerably, transitioning from relatively simple rule-based metrics toward more nuanced and fine-grained assessment frameworks. This evolution parallels the increasing sophistication of LLM capabilities and application scenarios. LLM-as-judge approaches have emerged as particularly promising, offering cost-effective evaluation solutions that can scale with the growing complexity of benchmarks. MT-bench [21] pioneered this approach while acknowledging inherent biases, subsequent studies have further elucidated critical issues.
Recent studies highlight several concerning limitations: Preference Leakage, as demonstrated by Li et al. (2025) [70], shows bias toward responses from models sharing architectural or training lineage with the judge model, compromising evaluation fairness. Contextual Sensitivity, revealed by Xu et al. (2025) [71], manifests as performance degradation when evaluating outputs dependent on external context, such as retrieval-augmented generation, where even state-of-the-art judges struggle with consistency. Reference Dependence is another issue, as evaluations often exhibit brittleness when reference solutions are unavailable or when multiple valid approaches exist, a common scenario in open-ended multi-turn interactions.
These challenges underscore the continued importance of human evaluation validation. Human-AI agreement checks are necessary to establish the reliability of automated evaluation methods, yet our analysis reveals that relatively few benchmarks incorporate substantial human agreement verification, representing a critical weakness in current evaluation frameworks, potentially allowing biases and inconsistencies to persist undetected.

### v5

@@ 2.1.1 Instruction Following Tasks in General
In addition to task-oriented evaluations, a complementary line of work has explored abstract reasoning benchmarks designed to test LLMs’ core reasoning capabilities in multi-turn settings. AQA-Bench (Yang et al., 2025b) evaluates LLMs in interactive environments requiring sequential decision-making, such as binary search and graph traversal (DFS, BFS). It emphasizes memory maintenance, procedural adherence, and planning over multiple turns, with both algorithmic and “embodied” (narrative) settings. WILT (Wason Inductive Logic Test) (Banatt et al., 2024) further assesses inductive reasoning by asking models to iteratively discover hidden rules through evidence gathering and hypothesis testing. These benchmarks are task-agnostic and explicitly designed to avoid memorization, targeting core skills such as logical consistency, strategic exploration, and hypothesis refinement.
@@
Besides these abstract reasoning evaluations, other studies have addressed multi-turn interactions from specialized task perspectives. For instance, WebLINX (Lù et al., 2024) tackles the problem of conversational web navigation, where an LLM-based agent must follow user instructions via dialogue to accomplish tasks on real websites. Tasks range from booking tickets to finding information, requiring the agent to understand natural language commands and manipulate a web page accordingly (click links, fill forms, etc.) over multiple turns. MULTITURNINSTRUCT (Han et al., 2025) proposes a systematic benchmark to probe LLMs’ ability to handle sequential, potentially conflicting instructions in a conversation. SysBench (Qin et al., 2025) is a benchmark specifically designed to evaluate how well LLMs adhere to system-level instructions (the hidden directives guiding an AI assistant) in multi-turn interactions. Emphasizing clear and explicit user intentions, we categorize two recent recommendation system works in this subsection: SAPIENT (Du et al., 2025) and ECR (Zhang et al., 2024b). SAPIENT (Du et al., 2025) is a multi-turn conversational recommender system that integrates a planning module for strategic dialogue management, while ECR (Zhang et al., 2024b) introduces an “empathetic”conversational recommender that enhances traditional recommendation dialogues with emotional awareness.
@@ 2.1.2 Instruction Following Tasks in Math
LLMs have demonstrated impressive performance in solving math problems via single-turn prompts, often generating detailed, step-by-step “chain-of-thought” solutions. For example, providing a few worked examples allows a 540B-parameter model to achieve state-of-the-art accuracy on the GSM8K math word problems (Wei et al., 2022). However, complex mathematical tasks frequently require LLMs to engage in multi-turn, instruction-following interactions that involve incremental reasoning, clarification questions, and iterative refinement based on interactive feedback (Liang et al., 2024; Romera-Paredes et al., 2024). In these instruction-following math tasks, users iteratively guide LLMs by providing clarifications, corrections, or additional contextual instructions. Such dynamic interactions not only enhance the models’ reasoning processes but also facilitate effective use of external computational tools, enabling the models to perform sophisticated tasks like simulating scenarios or executing complex calculations through code (Romera-Paredes et al., 2024; Lightman et al., 2023). Instruction-following tasks in math also encompass advanced skills such as follow-up questioning, errors diagnosing and correcting, and educational feedback delivering, thus capturing broader capabilities essential for deploying LLMs in diverse educational and problem-solving contexts.
Several approaches have leveraged multi-turn dialogue with LLMs to enhance mathematical reasoning. For example, Wu et al. (2024b) introduces MathChat-Agent, a framework where an LLM collaborates with a user-proxy agent (responsible for tool use, such as a Python solver) via iterative conversation. This approach solved competition-level math problems more effectively, improving accuracy by roughly 6% over standard single-turn chain-of-thought prompts. Similarly, Keating (2024) employs a multi-agent strategy in which two GPT-4 agents engage in debate-style interactions to reach a solution. This dual-agent zero-shot method achieved about 62.7% accuracy on the MATH benchmark—surpassing single-agent baselines and illustrating the benefits of peer deliberation in reasoning. Moving further, Xiong et al. (2025) propose a method to train LLMs to better combine tool usage with their own reasoning for complex tasks. It gathers trajectory-level feedback from users over multiple turns, to refine problem-solving strategies. The learning process is modeled as a Markov decision process (MDP) and adapts direct preference learning algorithms to multi-turn interactions that include external messages. In practice, the model is trained on multi-turn chats where a user first asks a question and then gives Python outputs in later turns. The accuracy of the method is evaluated on the GSM8K and MATH test sets.
@@
Several benchmarks now target multi-turn math dialogues. MathChat-Bench (Liang et al., 2024) extends GSM8K with four new tasks: follow-up QA, problem generation, error correction, and error analysis. The original problems are modified using GPT-4 to meet specific requirements. For instance, in the follow-up QA task, three rounds of dialogue are created: first, GSM8K test problems with ground-truth answers are presented; then GPT-4o generates two follow-up questions; finally, GPT-4 produces the final answers, which are verified or revised by two other LLMs and human annotators. LLMs are evaluated by both accuracy and scores, assigned by GPT-4, which measure their instruction follow-up abilities. Evaluations of state-of-the-art models on MathChat showed that while they excel at standard one-shot math questions, performance drops sharply on these multi-turn interactions requiring sustained reasoning and dialogue comprehension. To address this gap, the authors also released MathChatSync, a synthetic dialogue dataset for fine-tuning LLMs on conversational math problem solving. Fine-tuning on MathChatSync yielded notable improvements in multi-turn performance.
More broadly, the MINT benchmark (Wang et al., 2024f) evaluates multi-turn tool use and user feedback across domains (including math reasoning). MINT provides an automated framework where the LLM can call a Python interpreter and receive natural language feedback (simulated by GPT-4) in successive turns. Findings from MINT show that multi-turn tool-aided dialogues consistently improve problem-solving success (each tool use or feedback turn yields additional 1–17% accuracy gains). Interestingly, this evaluation also found that some models fine-tuned only on single-turn instructions (via standard supervised tuning or RLHF) underperform in multi-turn settings, suggesting that multi-turn-specific training is needed to excel in interactive math tasks.
Noticeably, several studies have found that LLMs struggle with generalizing to new problems. For example, Liang et al. (2024) show that math-specific LLMs lack adaptive behavior, and their difficulty with generating novel problems highlights their rigidity. Similarly, Macina et al. (2023) report that dialogue tutoring models do not generalize well to unseen math problems. To improve multi-turn math problem solving and generalization, many researchers propose using SFT (Liang et al., 2024; Macina et al., 2023; Wang et al., 2024f) or RLHF (Wang et al., 2024f; Xiong et al., 2025). Although RLHF can affect LLM–tool interactions and leveraging feedbacks (Wang et al., 2024f), studies consistently find that fine-tuning with preference data and instructions boosts downstream performance. Macina et al. (2023) demonstrate that small, fine-tuned models perform significantly better, in terms of correctness and equitable tutoring, than prompting a large model like ChatGPT.
@@ 2.1.3 Instruction Following Tasks in Coding
LLMs often struggle to produce correct code and perform self-debugging in a single pass, frequently requiring multi-turn or iterative interactions (Zhong et al., 2024; Chen et al., 2025e; Shi et al., 2024d). Instruction-following tasks in coding contexts commonly involve collaborative, iterative interactions, where users provide detailed instructions that the LLM translates into executable code. Through subsequent turns, the LLM iteratively integrates feedback, refines initial solutions, strategically plans modifications, executes and tests submodules, and performs debugging until satisfying the given specifications. Such iterative instruction-following tasks are essential to evaluate LLM performance and robustness in dynamic coding scenarios.
@@
Different frameworks and playgrounds are developed for evaluaing code generation qualities (Yang et al., 2023; Zheng et al., 2025). InterCode (Yang et al., 2023) introduces a generation pipeline for evaluating LLM coding quality through multi-turn interactions that simulate a real-world coding environment. The InterCode framework requires as input a natural language prompt paired with either an answer or a correct code block. The LLM is evaluated using one of three strategies: “Try-again”, where the execution output is fed back as an observation; ReACT, which terminates once the thought chain is complete; or Plan & Solve, which terminates when the plan is fully executed. The experiments involve three programming environments with Spider (Yu et al., 2018), MBPP (Austin et al., 2021) and NL2Bash (Lin et al., 2018) datasets. Zheng et al. (2025) propose a framework for systematically evaluating various prompting techniques for multi-turn code generation by LLMs. Their evaluation is conducted in a zero-shot setting using two competitive coding benchmarks, CodeContests (Li et al., 2022) and TACO (Li et al., 2023d). PyBench (Zhang et al., 2024c) proposes a unified benchmark to evaluate LLM Python coding ability in several categories such as chart analysis, software development, etc. For each task, LLMs are interacting with a code interpreter for a few turns before making a formal response. The LLMs are evaluated on the success rates and average turns to complete a unit test for each task.
@@
Toward effective code generations, CodeGEN (Nijkamp et al., 2023b) and CodeGEN2 (Nijkamp et al., 2023a) are a family of LLMs designed to generate programs from natural language descriptions in multiple turns. The models are evaluated using the Pass Rate metric on their Multi-Turn Programming Benchmark (MTPB), which comprises 115 expert-written problems. Each problem includes multi-step natural language prompts, created by human annotators who decompose the problem into sequential steps. Models are required to generate the complete solution from scratch. Chen et al. (2025d) propose to fine-tune existing LLMs with SFT and DPO. They introduce CodeSteer, a framework that guides LLMs through multiple rounds of interaction to generate code. In this system, the primary model (TaskLLM) produces responses, both in natural language and in code, while a supervisory agent (CodeSteerLLM) reviews these outputs using symbolic reasoning and self-answer checking to ensure correctness and provide refined guidance. They are fine-tuned and evaluated on subsets of SymBench, which comprises 37 symbolic tasks with adjustable complexity and includes a synthesized dataset of multi-round guidance/generation trajectories and guidance comparison pairs. OpenCodeInterpreter (Zheng et al., 2024b) proposes to fine-tune LLMs with CodeFeedback, a dataset of challenging LeetCode questions, incorporating multi-turn execution feedback (with code interpreters) and dialogues of human (synthesized by GPT-4). CodeAct (Wang et al., 2024e) collects an instruction-tuning dataset, CodeActInstruct, which contains 7,000 multi-turn interactions. PyInstruct (Zhang et al., 2024c) is used in PyBench for continuous pretraining and fine-tuning. Zheng et al. (2025) explore a wide range of prompting strategies for effective code generation, focusing on automatic re-prompting over multiple runs.
@@
SQL generation is a subcategory of coding generation that focuses more on data acquisition through large-scale data warehouses. Two benchmarks for SQL generation multi-turn LLMs are identified during the literature search. MMSQL (Guo et al., 2025) focuses on text-to-SQL generation tasks and introduces a Multi-type and Multi-turn text to-SQL test suite, which is a comprehensive benchmark engineered to evaluate the proficiency of LLMs in handling multi-turn text-to-SQL tasks across diverse question types. MMSQL contains a multi-agent framework anchored by a core Question Detector and Question Decomposer tasked with identifying question types and determining appropriate answering strategies. The framework includes two supportive agents: the Schema Selector, which identifies and provides the essential subset of a database schema, and the SQL Refiner, which is dedicated to refining SQL queries. EHRAgent (Shi et al., 2024a) demonstrates a specific application of SQL generation in healthcare settings, which speeds up the extraction and interaction of clinician information within electronic health record (EHR) systems. EHRAgent (Shi et al., 2024a) translates EHR question-answering into a tool-use planning process, which integrates query-specific medical information and formulates executable code plans through multi-turn dialogues. The model is evaluated based on their ability to reason across multiple tables and generate accurate, actionable insights from complex EHR data.
@@
For educational applications, TreeInstruct (Kargupta et al., 2024) transforms LLMs into Socratic instructor agents that guide users in debugging and writing better code. It operates via two roles: an instructor that generates tree-structured sequential questions, and a verifier that identifies tasks to help students understand, assess, and correct their code. To evaluate TreeInstruct, the authors introduce MULTIDEBUG, a dataset derived from LeetCode problems with expert-injected syntactic and conceptual bugs, assessed through both qualitative measures (relevance, indirectness, logical flow) and quantitative metrics (success rate).
Several compelling research questions recur in efforts to adapt LLMs for reliable code generation in multi-turn settings. First, when explicit cues are unavailable, how should an LLM decide between purely textual reasoning and executing programmatic actions to reach a solution (Chen et al., 2025e)? Second, which iterative interaction protocols best support incremental refinement—enabling the model to diagnose errors, incorporate feedback, and converge on a final answer (Yang et al., 2023; Zheng et al., 2025; Nijkamp et al., 2023b; a; Wang et al., 2024e)? Third, how should we evaluate coding performance across programming languages, task types, and application domains in a way that is both comparable and representative of real-world use (Yang et al., 2023; Zhang et al., 2024c)?
@@ 2.1.4 Discussions
The evaluation landscape has diversified considerably, transitioning from relatively simple rule-based metrics toward more nuanced and fine-grained assessment frameworks. This evolution parallels the increasing sophistication of LLM capabilities and application scenarios. LLM-as-judge approaches have emerged as particularly promising, offering cost-effective evaluation solutions that can scale with the growing complexity of benchmarks. MT-bench (Zheng et al., 2023) pioneers this approach while acknowledging inherent biases, subsequent studies have further elucidated critical issues.
Recent studies highlight several concerning limitations: Preference Leakage, as demonstrated by Li et al. (2026), shows bias toward responses from models sharing architectural or training lineage with the judge model, compromising evaluation fairness. Contextual Sensitivity, revealed by Xu et al. (2025), manifests as performance degradation when evaluating outputs dependent on external context, such as retrieval-augmented generation, where even state-of-the-art judges struggle with consistency. Reference Dependence is another issue, as evaluations often exhibit brittleness when reference solutions are unavailable or when multiple valid approaches exist, a common scenario in open-ended multi-turn interactions.
These challenges underscore the continued importance of human evaluation validation. Human-AI agreement checks are essential for establishing the reliability of automated evaluation methods, yet relatively few benchmarks incorporate substantial human agreement verification—a critical gap that risks allowing biases and inconsistencies to persist undetected.